OpenStudy (anonymous):

Suppose the population of a country in 1985 was 145 million. In 1995 it was 190 million. Use the formula P = Ae* kt to setup an expression to estimate the population of the country in 2005.

OpenStudy (campbell_st):

is t the time in years since 1985..?

OpenStudy (anonymous):

i think so

OpenStudy (campbell_st):

ok... so that means your initial population A = 145 at t = 10 the P = 190 so your equation becomes \[190 = 145e^{10k}\] you need to find the growth constant k so \[\frac{190}{145} = e^{10k}\] take the ln of both sides to find k \[\ln(\frac{190}{145}) = 10k\] when you know k... substitute it into the equation \[P = 145e^{kt}\] to find the population in 2005... let t = 20.. hope that makes sense...

OpenStudy (anonymous):

yep thanks